find the coefficient of 1/y2 in (y+c3/y2)10

Grade:11

3 Answers

pratish shrivastava
13 years ago

Dear kaif ali,

we know that,         Tr=nCr(I)n-r(II)r

=>T4=10C4(y)10-4(c3/y2)4

=>T4=210(y)6(c12/y8)

=>T4=210(c12/y2)

so, coefficient of  i/y2  is  210c12

I hope you understand 

 

Gurwinder Kaur
13 years ago

(y+c3/y2)10  = (10y+10c3)/y2

=1/y2*(10y+10c3)

coeff. of 1/y2 is (10y+10c3)

approve if u  like

Karteek Ravindra Bhat
13 years ago

w.k.t  

 

Tr+1=10Cr.y10-r[(c3)/y2]r

      therefore by simplifyin we get

y2r-10+r=y2

therefore 2r-10+r=2

so r=4

simplify in eqn 1

v get t5=10c4.y2.c12

so coeff of y2 =210*c12

!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!